OpenStudy (anonymous):

Using point-slope form, write the equation of the line that passes through (2, –1) and has a slope of –5. A. (y – 1) = –5(x + 2) B. (y – 2) = –5(x – 1) C. (y + 1) = –5(x – 2) D. (y + 2) = –5(x – 1)

OpenStudy (anonymous):

y - y1 = m(x - x1) <== point slope formula slope(m) = -5 (2,-1)...x1 = 2 and y1 = -1 now we sub y - (-1) = -5(x - 2) y + 1 = -5(x - 2) <=== point slope form

OpenStudy (anonymous):

I think it's D but I'm not entirely sure. Is it C or D?

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

no problem :)